Encircle the odd one of the following 
(a) (–3, 3) (b) (–5, 5) (c) (–6, 1) (d) (–8, 8) 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (c) is correct here
–3 + 3 = 0
–5 + 5 = 0
–8 + 8 = 0
–6 + 1 = -5
Hence, (–6, 1) is the odd one.

Thanks